# Gorenstein singularities with $\mathbb{G}_{m}$-action and moduli spaces of holomorphic differentials

### v1

@@ 5.13. Nonvarying strata without obstructions
Consider the nonvarying stratum $\mathbb{P}\mathcal{H}(2,1,1)$. The defining equations of the corresponding singularity $(Y,q)$ have been given in Section 5.3, where $(Y,q)$ is not a local complete intersection. Nevertheless, with the defining equations at hand, we can compute its versal deformation space by using the software Singular as follows:
  • LIB "sing.lib";
  • ring R = 0,(x,y,z),dp;
  • ideal I = xy, yz2-y4, xz2-x3z, z2-y3-x2z;
  • deform(I);
$\begin{pmatrix}xy\\ yz^{2}-y^{4}\\ xz^{2}-x^{3}z\\ z^{2}-y^{3}-x^{2}z\\ \end{pmatrix}$ $+\lambda_{1}\begin{pmatrix}0\\ yz\\ xz\\ z\\ \end{pmatrix}+\lambda_{2}\begin{pmatrix}0\\ y^{2}\\ xy\\ y\\ \end{pmatrix}+\lambda_{3}\begin{pmatrix}0\\ xy\\ x^{2}\\ x\\ \end{pmatrix}+\lambda_{4}\begin{pmatrix}0\\ y\\ x\\ 1\\ \end{pmatrix}$
$+\lambda_{5}\begin{pmatrix}z\\ xz^{2}\\ y^{2}z\\ 0\\ \end{pmatrix}+\lambda_{6}\begin{pmatrix}y\\ xyz\\ y^{3}\\ 0\\ \end{pmatrix}+\lambda_{7}\begin{pmatrix}x\\ x^{2}z\\ xy^{2}\\ 0\\ \end{pmatrix}+\lambda_{8}\begin{pmatrix}1\\ xz\\ y^{2}\\ 0\\ \end{pmatrix}.$
The weights of $x$, $y$, and $z$ are $3$, $4$, and $6$, respectively. The weights of the deformation parameters $(\lambda_{1},\ldots,\lambda_{8})$ can be read off from those of $(z,y^{2},xz,y^{3},xy/z,x,y,xy)$ which are $(6,8,9,12,1,3,4,7)$. Therefore, $\mathbb{P}\mathcal{H}(2,1,1)$ can be compactified by the versal deformation space of $(Y,q)$ as the weighted projective space
${\rm Def}(Y,q)\cong\mathbb{P}(1,3,4,6,7,8,9,12).$
Remark 5.3.
The structure of $\mathbb{P}{\rm Def}(Y,q)^{-}$ for monomial singularities with fixed semigroups has also been understood fairly well in low genus; see [Nak08, Ste23].
Moreover, as mentioned in Section 5.12, we can add entries of 00 into the signature $\mu$, i.e., zeros of order 00, to generalize the setting of nonvarying strata. Nevertheless, increasing the number of zero entries can lead to more complicated deformations for the corresponding singularities. For example, if $(Y,q)$ is the elliptic $n$-fold point, then $T^{2}$ is nonzero for $n\geq 6$ and $\mathbb{P}{\rm Def}^{-}(Y,q)$ is not a weighted projective space for $n=6$; see [Ste25, Proposition 1.2] and [LP19, Theorem 1.5.7].
@@ 6. Singularities with bounded $\alpha$-invariants
$\chi_{1}^{\log}=\sum_{\lambda=1}^{\ell}h^{0}(C,D_{\lambda})\geq\frac{1}{4}(2g-2+n)\ell.$ (12)
@@
$\sum_{\lambda=1}^{\ell}\ell_{\lambda,i}=\frac{1}{2}(m_{i}+2)\ell\quad{\rm and}\quad\sum_{\lambda=1}^{\ell}\sum_{i=1}^{n}\ell_{\lambda,i}=(g-1+n)\ell.$ (13)
@@
$\chi_{1}^{\log}$ $\leq\sum_{\lambda=1}^{\ell}\left(1+\frac{1}{2}\left(2g-2+n-\sum_{i=1}^{n}l_{\lambda,i}\right)\right)=\frac{1}{2}(g+1)\ell.$ (14)
@@
Additionally, since adding an ordinary marked point (i.e., a zero of order 00) does not change the value of $\chi_{1}^{\log}$, we can first classify the cases where every entry in $\mu$ is positive, and then add zero entries into $\mu$ up to the bound given by (12).
@@ 6.1. The case $n=4$
Next, consider the case $m_{1}\geq 2$. Then Clifford’s bound holds for some $D_{\lambda}$ where $20<\deg D_{\lambda}<2g-$ 2, and hence the underlying curve $C$ is hyperelliptic. If a zero $z_{i}$ is a Weierstrass point, the divisor class $D_{a_{i}+1}$ is of the form $K_{C}-z_{i}-E_{i}$, where $E_{i}$ is an effective divisor not containing $z_{i}$. Then $D_{a_{i}+1}$ does not achieve the maximum of Clifford’s bound, leading to a contradiction. Therefore, the remaining possibility is that the four zeros form two hyperelliptic conjugacy pairs, where $m_{1}=m_{2}=a$ and $m_{3}=m_{4}=b$ (up to reordering). In this case, (12) attains equality.
@@ 6.3. The nonhyperelliptic case with $n=3$
We first make the following observation that will be used frequently for all nonhyperelliptic cases. If the underlying curve $C$ is nonhyperelliptic, then for any divisor class $D$ with $20<\deg D<2g-$ 2, we have
$h^{0}(C,D)\leq\begin{cases}\frac{1}{2}(\deg D+1)&{\rm if}\ \deg D\ {\rm is\ odd},\\ \frac{1}{2}\deg D&{\rm if}\ \deg D\ {\rm is\ even}.\end{cases}$ (15)
@@
$\chi_{1}^{\log}$ $\leq\sum_{\lambda=1}^{\ell}\frac{1}{2}\left(2g-1+n-\sum_{i=1}^{n}l_{\lambda,i}\right)-\frac{1}{2}N^{+}+a_{1}$ (16)
@@
For the case $n=3$, we assume that $m_{1}\geq m_{2}\geq m_{3}$ as before, and then $a_{1}\leq a_{2}\leq a_{3}$. Comparing (16) with (12), it follows that $2N^{+}+\ell\leq 4a_{1}$, and hence $m_{1}\leq 3$. It is straightforward to check the remaining possibilities $\mu=(3,3,2)$, $(3,2,1)$, $(2,2,2)$, and $(2,1,1)$ explicitly. Only $\mu=(2,1,1)$ and $(2,2,2)^{\operatorname{even}}$ work, whose $\alpha$-invariants have been computed in Sections 5.3 and 5.11, respectively. Additionally, if we add an ordinary marked point to them, by Section 5.12, it would make their $\alpha$-invariants less than $3/8$.
@@ 6.4. The nonhyperelliptic case with $n=2$
Consider first the case $(g-1,g-1)$, where $\ell=g$ and $a_{1}=a_{2}=1$. Then $N^{+}=g-2\leq 2$ and $g\leq 4$. The cases are $\mu=(2,2)^{\operatorname{odd}}$ and $(3,3)^{\operatorname{nonhyp}}$, whose $\alpha$-invariants have been computed in Sections 5.2 and 5.6, respectively. In both cases, no ordinary marked points can be added to them.
@@
If $d=4$, then $k_{2}=1$ and $k_{1}=2$, which is the case $\mu=(7,3)$ and the nonhyperelliptic Clifford’s bound in (15) attains its maximum for all divisor classes $D_{\lambda}$ that appear in the range $20<\deg D_{\lambda}<2g-$ 2. This is equivalent to $h^{0}(C,2p_{1}+p_{2})=2$ (and $C$ is not hyperelliptic by the assumption). Using the method in Section 5, we find that $\mathfrak{m}/\mathfrak{m}^{2}\pmod{\widetilde{\mathfrak{m}}^{9}}$ is generated by $x=t_{1}^{2}\oplus t_{2}$ and $y=t_{1}^{5}\oplus 0$, and the defining equation of the singularity is $y(y^{2}-x^{5})$. Alternatively, the condition $h^{0}(C,2p_{1}+p_{2})=2$ implies that $h^{0}(C,5p_{1}+p_{2})=3$. We can take a plane sextic defined by $y^{3}f(x,y)-x^{5}$, where $f(x,y)$ is a general polynomial of degree three. The line $y=0$ cuts out a divisor of type $5p_{1}+p_{2}$, where $p_{1}$ is at the origin and $p_{2}$ is at infinity. Moreover, the projection from $p_{1}$ induces a $g^{1}_{3}$ with a fiber given by $2p_{1}+p_{2}$. Then the normalization of this plane sextic gives an element in $\mathbb{P}\mathcal{H}(7,3)$ which further satisfies $h^{0}(C,2p_{1}+p_{2})=2$.
@@ 6.5. The nonhyperelliptic case with $n=1$
$\sum_{i=1}^{g}k_{i}\leq g^{2}-1.$ (17)
By the definition of $H_{p}$, for $i\geq 1$, we have $h^{0}(C,(k_{i}-1)p)=i-1$ and $h^{0}(C,k_{i}p)=i$. Since we have assumed that the underlying curve $C$ is not hyperelliptic, by Clifford’s theorem, we have $h^{0}(C,k_{i}p)=i\leq(k_{i}+1)/2$, which implies that $k_{i}\geq 2i-1$ for $12\leq i\leq g-$ 1. It follows that
$\sum_{i=1}^{g}k_{i}\geq\sum_{i=2}^{g-1}(2i-1)+(2g-2)=g^{2}-2.$ (18)
@@
First, consider the case that $\sum_{i=1}^{g}k_{i}=g^{2}-2$. Then $H_{p}=\{0,3,5,\ldots,2g-3,2g-2,2g,2g+1,\ldots\}$. Since $H_{p}$ is a semigroup, it implies that $26=3+3\geq 2g-$ 2, and hence $g\leq 4$.
@@
If $k_{2}=3\in H_{p}$ and if $46\leq 2g-$ 4 (i.e., $g\geq 5$), then $H_{p}=\{0,3,6,7,9,\ldots,2g-5,2g-3,2g-2,2g,2g+1,\ldots\}$. It implies that $210=3+7\geq 2g-$ 2 and $g\leq 6$. In particular, for $g=6$, the only possible case is $H_{p}=\{0,3,6,7,9,10,12,13,\ldots\}$ generated by $3$ and $7$, where $G_{p}=\{1,2,4,5,8,11\}$. In this case, the defining equation of the singularity is $y^{3}-x^{7}$. We remark that it was verified in [Ste23, Table 1] that the locus of smooth pointed curves with this semigroup is non-empty.
If $k_{2}=4\in H_{p}$, then $H_{p}=\{0,4,5,7,\ldots,2g-5,2g-3,2g-2,2g,2g+1,\ldots\}$. It implies that $28=4+4\geq 2g-$ 2 and $g\leq 5$.
@@ 7. Slopes of singularities and effective divisors
Motivated by [AFS16, (1.2) and Theorem 4.2], given a singularity class $(Y,q)\in S(\mu)$ for a partition $\mu=(m_{1},\ldots,m_{n})$ of $22g-$ 2, we define the slope of $(Y,q)$ to be
$\begin{split}s(Y,q)&\coloneqq\frac{13\chi_{1}(Y)-\chi_{2}(Y)}{\chi_{1}(Y)}\\ &=12-\frac{(2g-2+n-\sum_{i=1}^{n}\frac{1}{m_{i}+1})\ell}{\chi_{1}(Y)},\end{split}$ (19)
@@
The above definition mimics the slope for a one-parameter family of stable curves $B$ in ¯$\overline{\mathcal{M}}_{g}$, which is defined by $s(B)=\deg_{B}\delta/\deg_{B}\lambda_{1}$, as well as the slope for a divisor class $D=a\lambda_{1}-b\delta$ in ¯$\overline{\mathcal{M}}_{g}$, which is defined by $s(D)=a/b$; see [HM90]. Additionally, by Theorem 2.6 (ii), smooth pointed curves $(C,p_{1},\ldots,p_{n})\in\mathbb{P}{\rm Def}^{-}_{s}(Y,q)$ have the same graded quotients of Rees filtrations, which determine $\chi_{m}(Y)$ completely for all $m$. Therefore, we also define the slope of the canonical divisor $\sum_{i=1}^{n}m_{i}p_{i}$ in $C$ to be $s(Y,q)$.
@@
We remark that if $(Y,q)\in S(2^{g-1})^{\operatorname{odd}}$ is from a special choice such that $h^{0}(C,p_{1}+\cdots+p_{g-1})\geq 3$, then the value of $s(Y,q)$ can be different from the above.
@@
We demonstrate for the case when $g$ is odd. In this case, we have $\ell=g+1$, $a_{1}=1$, and $a_{i}=(g+1)/2$ for $12\leq i\leq g-$ 1. In particular,
@@
It is a long standing question to determine the lower bound for the slopes of effective divisors in ¯$\overline{\mathcal{M}}_{g}$. Harris–Morrison conjectured that the Brill–Noether divisors have the lowest slope given by $6+12/(g+1)$; see [HM98, Conjecture (6.63)]. Although starting from $g=10$ a series of counterexamples have been discovered, their slopes are still close to the Brill–Noether slope; see [Tan98, FP05, Far09]. In particular, no effective divisors with slope $\leq 6$ have been found. Conversely, one can construct moving curves in ¯$\overline{\mathcal{M}}_{g}$ to bound the slopes of effective divisors by using their intersection numbers, where a moving curve passes through a general element in ¯$\overline{\mathcal{M}}_{g}$. Nevertheless, all known lower bounds constructed this way tend to zero as $g$ approaches infinity; see [Fed08, Che11, Pan12]. We refer to [CFM13] for a survey about this fantastic question.
Under the same assumption as in [AFS16, Theorem 4.2], we will illustrate an amusing approach to bound the slopes of effective divisors in ¯$\overline{\mathcal{M}}_{g}$ by using the slopes of $(Y,q)\in S(\mu)$ defined above. The following statement is an elaboration of Theorem 1.5.
@@
Let $\mu=(m_{1},\ldots,m_{n})$ be a partition of $22g-$ 2 with $n\geq g-1$ such that $\mu\neq(2^{g-1})^{\operatorname{even}}$. Let $(C,p_{1},\ldots,p_{n})\in\mathbb{P}\mathcal{H}(\mu)$ be a general element and let $(Y,q)\in S(\mu)$ be the corresponding singularity class. Suppose the discriminant locus of singular curves in ${\rm Def}^{-}(Y,q)$ is a $\mathbb{Q}$-Cartier divisor and the locus of worse-than-nodal deformations of $(Y,q)$ has codimension at least $2$. Then $s(Y,q)$ is a lower bound for the slopes of effective divisors in ¯$\overline{\mathcal{M}}_{g}$.

### v2

@@ 5.15. Nonvarying hyperelliptic loci
Suppose $2a_{i}+1$ and $b_{j}+1$ are divisible by an integer $d\geq 4$ for all $i$ and $j$. Let $k=2\ell/d$. Then
$kC_{\infty}=\sum_{i=1}^{m}\frac{2(2a_{i}+1)}{d}w_{i}+\sum_{j=1}^{n}\frac{2(b_{j}+1)}{d}(u_{i}+v_{i})$
is a Cartier divisor. Since $d\geq 4$, we have the even coefficient of $w_{i}$ equal to $2(2a_{i}+1)/d<2a_{i}$, and hence $kC_{\infty}$ is strictly less effective than the canonical divisor $\sum_{i=1}^{m}2a_{i}w_{i}+\sum_{j=1}^{n}b_{j}(u_{i}+v_{i})$, which implies that $\deg kC_{\infty}<2g-2$. Therefore, we can apply Proposition 5.7 to conclude that in this case all deformations in $\mathbb{P}{\rm Def}^{-}_{s}(Y,q)$ are hyperelliptic.
Example 5.9.
Without loss of generality, we reorder the zeros such that $a_{1}\geq\cdots\geq a_{m}$. Moreover, suppose $a_{1}\geq 2$ and $2a_{1}+1>2(2a_{i}+1)$ for all $a_{i}\neq a_{1}$. Let $k=2\ell/(2a_{1}+1)$. Then we have
$\lfloor kC_{\infty}\rfloor=\sum_{i=1}^{m}\left\lfloor\frac{2(2a_{i}+1)}{2a_{1}+1}\right\rfloor w_{i}+\sum_{j=1}^{n}\left\lfloor\frac{2(b_{j}+1)}{2a_{1}+1}\right\rfloor(u_{j}+v_{j}).$
Note that the coefficient of $w_{i}$ is $2$ if $a_{i}=a_{1}$ and is 0 if $a_{i}\neq a_{1}$. Since $2a_{1}>2$, the divisor $\lfloor kC_{\infty}\rfloor$ is strictly less effective than the corresponding canonical divisor. Therefore, we can apply Proposition 5.7 to conclude that in this case all deformations in $\mathbb{P}{\rm Def}^{-}_{s}(Y,q)$ are hyperelliptic.
@@ 6. Singularities with bounded $\alpha$-invariants
(12) $\chi_{1}^{\log}=\sum_{\lambda=1}^{\ell}h^{0}(C,D_{\lambda})\geq\frac{1}{4}(2g-2+n)\ell.$
@@
(13) $\sum_{\lambda=1}^{\ell}\ell_{\lambda,i}=\frac{1}{2}(m_{i}+2)\ell\quad{\rm and}\quad\sum_{\lambda=1}^{\ell}\sum_{i=1}^{n}\ell_{\lambda,i}=(g-1+n)\ell.$
@@
(14) $\chi_{1}^{\log}$ $\leq\sum_{\lambda=1}^{\ell}\left(1+\frac{1}{2}\left(2g-2+n-\sum_{i=1}^{n}l_{\lambda,i}\right)\right)=\frac{1}{2}(g+1)\ell.$
@@
Additionally, since adding an ordinary marked point (i.e., a zero of order 0) does not change the value of $\chi_{1}^{\log}$, we can first classify the cases where every entry in $\mu$ is positive, and then add zero entries into $\mu$ up to the bound given by (12).
@@ 6.1. The case $n=4$
Next, consider the case $m_{1}\geq 2$. Then Clifford’s bound holds for some $D_{\lambda}$ where $0<\deg D_{\lambda}<2g-2$, and hence the underlying curve $C$ is hyperelliptic. If a zero $z_{i}$ is a Weierstrass point, the divisor class $D_{a_{i}+1}$ is of the form $K_{C}-z_{i}-E_{i}$, where $E_{i}$ is an effective divisor not containing $z_{i}$. Then $D_{a_{i}+1}$ does not achieve the maximum of Clifford’s bound, leading to a contradiction. Therefore, the remaining possibility is that the four zeros form two hyperelliptic conjugacy pairs, where $m_{1}=m_{2}=a$ and $m_{3}=m_{4}=b$ (up to reordering). In this case, (12) attains equality.
@@ 6.3. The nonhyperelliptic case with $n=3$
We first make the following observation that will be used frequently for all nonhyperelliptic cases. If the underlying curve $C$ is nonhyperelliptic, then for any divisor class $D$ with $0<\deg D<2g-2$, we have
(15) $h^{0}(C,D)\leq\begin{cases}\frac{1}{2}(\deg D+1)&{\rm if}\ \deg D\ {\rm is\ odd},\\ \frac{1}{2}\deg D&{\rm if}\ \deg D\ {\rm is\ even}.\end{cases}$
@@
(16) $\chi_{1}^{\log}$ $\leq\sum_{\lambda=1}^{\ell}\frac{1}{2}\left(2g-1+n-\sum_{i=1}^{n}l_{\lambda,i}\right)-\frac{1}{2}N^{+}+a_{1}$
@@
For the case $n=3$, we assume that $m_{1}\geq m_{2}\geq m_{3}$ as before, and then $a_{1}\leq a_{2}\leq a_{3}$. Comparing (16) with (12), it follows that $2N^{+}+\ell\leq 4a_{1}$, and hence $m_{1}\leq 3$. It is straightforward to check the remaining possibilities $\mu=(3,3,2)$, $(3,2,1)$, $(2,2,2)$, and $(2,1,1)$ explicitly. Only $\mu=(2,1,1)$ and $(2,2,2)^{\operatorname{even}}$ work, whose $\alpha$-invariants have been computed in Sections 5.3 and 5.12, respectively. Additionally, if we add an ordinary marked point to them, by Section 5.14, it would make their $\alpha$-invariants less than $3/8$.
@@ 6.4. The nonhyperelliptic case with $n=2$
Consider first the case (g1,g1)(g-1,g-1), where $\ell=g$ and $a_{1}=a_{2}=1$. Then $N^{+}=g-2\leq 2$ and $g\leq 4$. The cases are $\mu=(2,2)^{\operatorname{odd}}$ and $(3,3)^{\operatorname{nonhyp}}$, whose $\alpha$-invariants have been computed in Sections 5.2 and 5.6, respectively. In both cases, no ordinary marked points can be added to them.
@@
If $d=4$, then $k_{2}=1$ and $k_{1}=2$, which is the case $\mu=(7,3)$ and the nonhyperelliptic Clifford’s bound in (15) attains its maximum for all divisor classes $D_{\lambda}$ that appear in the range $0<\deg D_{\lambda}<2g-2$. This is equivalent to $h^{0}(C,2p_{1}+p_{2})=2$ (and $C$ is not hyperelliptic by the assumption). Using the method in Section 5, we find that $\mathfrak{m}/\mathfrak{m}^{2}\pmod{\widetilde{\mathfrak{m}}^{9}}$ is generated by $x=t_{1}^{2}\oplus t_{2}$ and $y=t_{1}^{5}\oplus 0$, and the defining equation of the singularity is $y(y^{2}-x^{5})$. Alternatively, the condition $h^{0}(C,2p_{1}+p_{2})=2$ implies that $h^{0}(C,5p_{1}+p_{2})=3$. We can take a plane sextic defined by $y^{3}f(x,y)-x^{5}$, where $f(x,y)$ is a general polynomial of degree three. The line $y=0$ cuts out a divisor of type $5p_{1}+p_{2}$, where $p_{1}$ is at the origin and $p_{2}$ is at infinity. Moreover, the projection from $p_{1}$ induces a $g^{1}_{3}$ with a fiber given by $2p_{1}+p_{2}$. Then the normalization of this plane sextic gives an element in $\mathbb{P}\mathcal{H}(7,3)$ which further satisfies $h^{0}(C,2p_{1}+p_{2})=2$.
@@ 6.5. The nonhyperelliptic case with $n=1$
(17) $\sum_{i=1}^{g}k_{i}\leq g^{2}-1.$
By the definition of $H_{p}$, for $i\geq 1$, we have $h^{0}(C,(k_{i}-1)p)=i-1$ and $h^{0}(C,k_{i}p)=i$. Since we have assumed that the underlying curve $C$ is not hyperelliptic, by Clifford’s theorem, we have $h^{0}(C,k_{i}p)=i\leq(k_{i}+1)/2$, which implies that $k_{i}\geq 2i-1$ for $2\leq i\leq g-1$. It follows that
(18) $\sum_{i=1}^{g}k_{i}\geq\sum_{i=2}^{g-1}(2i-1)+(2g-2)=g^{2}-2.$
@@
First, consider the case that $\sum_{i=1}^{g}k_{i}=g^{2}-2$. Then $H_{p}=\{0,3,5,\ldots,2g-3,2g-2,2g,2g+1,\ldots\}$. Since $H_{p}$ is a semigroup, it implies that $6=3+3\geq 2g-2$, and hence $g\leq 4$.
@@
If $k_{2}=3\in H_{p}$ and if $6\leq 2g-4$ (i.e., $g\geq 5$), then $H_{p}=\{0,3,6,7,9,\ldots,2g-5,2g-3,2g-2,2g,2g+1,\ldots\}$. It implies that $10=3+7\geq 2g-2$ and $g\leq 6$. In particular, for $g=6$, the only possible case is $H_{p}=\{0,3,6,7,9,10,12,13,\ldots\}$ generated by $3$ and $7$, where $G_{p}=\{1,2,4,5,8,11\}$. In this case, the defining equation of the singularity is $y^{3}-x^{7}$. We remark that it was verified in [Ste23, Table 1] that the locus of smooth pointed curves with this semigroup is non-empty.
If $k_{2}=4\in H_{p}$, then $H_{p}=\{0,4,5,7,\ldots,2g-5,2g-3,2g-2,2g,2g+1,\ldots\}$. It implies that $8=4+4\geq 2g-2$ and $g\leq 5$.
@@ 7. Slopes of singularities and effective divisors
Motivated by [AFS16, (1.2) and Theorem 4.2], given a singularity class $(Y,q)\in S(\mu)$ for a partition $\mu=(m_{1},\ldots,m_{n})$ of 2g22g-2, we define the slope of $(Y,q)$ to be
(19) $\begin{split}s(Y,q)&\coloneqq\frac{13\chi_{1}(Y)-\chi_{2}(Y)}{\chi_{1}(Y)}\\ &=12-\frac{(2g-2+n-\sum_{i=1}^{n}\frac{1}{m_{i}+1})\ell}{\chi_{1}(Y)},\end{split}$
@@
The above definition mimics the slope for a one-parameter family of stable curves $B$ in $\overline{\mathcal{M}}_{g}$, which is defined by $s(B)=\deg_{B}\delta/\deg_{B}\lambda_{1}$, as well as the slope for a divisor class $D=a\lambda_{1}-b\delta$ in $\overline{\mathcal{M}}_{g}$, which is defined by $s(D)=a/b$; see [HM90]. Additionally, by Theorem 2.6 (ii), smooth pointed curves $(C,p_{1},\ldots,p_{n})\in\mathbb{P}{\rm Def}^{-}_{s}(Y,q)$ have the same graded quotients of Rees filtrations, which determine $\chi_{m}(Y)$ completely for all $m$. Therefore, we also define the slope of the canonical divisor $\sum_{i=1}^{n}m_{i}p_{i}$ in $C$ to be $s(Y,q)$.
@@
In this case, we can apply the method in Section 5.8 to describe the generators of $\mathfrak{m}/\mathfrak{m}^{2}\pmod{\tilde{\mathfrak{m}}^{4}}$ for the singularity $(Y,q)$ as follows:
$x_{i}$ $=0\oplus\cdots\oplus 0\oplus t_{i}^{2}\oplus 0\oplus\cdots\oplus 0,\quad i=1,\ldots,g-1,$
$y_{j}$ $=0\oplus\cdots\oplus 0\oplus t_{j}^{3}\oplus 0\oplus\cdots\oplus 0\oplus t_{g-1}^{3},\quad j=1,\ldots,g-2.$
We remark that if $(Y,q)\in S(2^{g-1})^{\operatorname{odd}}$ is from a special choice such that $h^{0}(C,p_{1}+\cdots+p_{g-1})\geq 3$, then the value of $s(Y,q)$ and the structure of the singularity can be different from the above generic case.
@@
We demonstrate for the case when $g$ is odd. In this case, we have $\ell=g+1$, $a_{1}=1$, and $a_{i}=(g+1)/2$ for $2\leq i\leq g-1$. In particular,
@@
It is a long standing question to determine the lower bound for the slopes of effective divisors in $\overline{\mathcal{M}}_{g}$. Harris–Morrison conjectured that the Brill–Noether divisors have the lowest slope given by $6+12/(g+1)$; see [HM98, Conjecture (6.63)]. Although starting from $g=10$ a series of counterexamples have been discovered, their slopes are still close to the Brill–Noether slope; see [Tan98, FP05, Far09]. In particular, no effective divisors with slope $\leq 6$ have been found. Conversely, one can construct moving curves in $\overline{\mathcal{M}}_{g}$ to bound the slopes of effective divisors by using their intersection numbers, where a moving curve passes through a general element in $\overline{\mathcal{M}}_{g}$. Nevertheless, all known lower bounds constructed this way tend to zero as $g$ approaches infinity; see [Fed08, Che11, Pan12]. We refer to [CFM13] for a survey about this fantastic question. Under the same assumption as in [AFS16, Theorem 4.2], we will illustrate an amusing approach to bound the slopes of effective divisors in $\overline{\mathcal{M}}_{g}$ by using the slopes of $(Y,q)\in S(\mu)$ defined above. The following statement is an elaboration of Theorem 1.6.
@@
Let $\mu=(m_{1},\ldots,m_{n})$ be a partition of 2g22g-2 with $n\geq g-1$ such that $\mu\neq(2^{g-1})^{\operatorname{even}}$. Let $(C,p_{1},\ldots,p_{n})\in\mathbb{P}\mathcal{H}(\mu)$ be a general element and let $(Y,q)\in S(\mu)$ be the corresponding singularity class. Suppose the discriminant locus of singular curves in ${\rm Def}^{-}(Y,q)$ is a $\mathbb{Q}$-Cartier divisor and the locus of worse-than-nodal deformations of $(Y,q)$ has codimension at least $2$. Then $s(Y,q)$ is a lower bound for the slopes of effective divisors in $\overline{\mathcal{M}}_{g}$.